# More About the Spontaneous Breaking of Time Reversal in de Sitter Space

### v1

@@
It is widely thought that the quantum theory of de Sitter space requires the existence of a physical observer in the static patch. What exactly is meant by an observer is unclear; it could be anything from a few photons with energy just above the Gibbons-Hawking temperature to a gravitationally bound cluster of galaxies. In a recent note I explained that the need for observers can arise from the spontaneous breaking of time-reversal symmetry. This longer paper expands on the subject, filling in conceptual arguments that were implicit but not explicitly stated in the earlier paper.
@@ 1 Introduction
There is a widespread belief that the quantum mechanics of de Sitter space only makes sense if there is an observer present [1]. Discrepancies between semiclassical gravity and static patch holography have been resolved by introducing a feature localized at the pode of the static patch. According to the terminology of [1][2][3][4] this feature is called an “observer.” I put “observer” in quotes because I don’t believe it necessarily has much to do with observation by almost classical apparatuses. “Observers” comprise a more general class than observers. Observers, among other requirements, are heavy enough to be localized near the pode; observers can be much lighter and can have wavelengths of cosmic scale.
@@
The ability of an observer to resolve the discrepancies is surprising because the semiclassical calculation may involve bulk degrees of freedom localized extremely far from the pode—for example near the horizon—and yet no matter how far away the pode is from the region of the semiclassical calculation, the presence of an observer at the pode changes things enough to resolve the discrepancy. On the face of it this is a form of non-locality forbidden in bulk quantum field theory.
One thing about “observers” is that they should have some form of time-reversal asymmetry to distinguish future and past. I’ll call such an asymmetric feature a “clock.”
@@
Generally the literature on this subject does not spell out exactly what is meant by an observer. In fact it seems that it could be anything from a galaxy to a tiny out-of-equilibrium feature in the Hawking radiation—no more than a handful of photons with energy of order the Gibbons-Hawking temperature. Calling such a feature an observer seems misleading to me; I prefer the terms feature, lump, or fluctuation to the overly suggestive observer but to avoid the introduction of new terminology I’ll stick with “observer.” But it is important to keep in mind that “observer” is just a name for an feature that could be almost anything.
To make matters more extreme, not only can the observer be extremely far from the region of the semiclassical calculation; it may also be a very feeble perturbation and still resolve the mismatch. It can be any “observer” as long as it can be localized in the bulk, far from the horizon. This means that its wavelength should be a small fraction of the de Sitter length $\ell.$
@@
Now that I’ve explained the “observer” let’s turn to the region $h$ in figure 1. I’ll refer to it as the laboratory (without quotes). It is the region of interest where experiments take place, the experiments being described by a set of correlation functions localized near $h.$
@@
In a recent note [5] I argued that the need for an “observer” has nothing to do with the consistency of quantum mechanics; it is due to a spontaneous symmetry breaking (SSB) of time-reversal symmetry. The “observer” is just the small feature which kicks off an instability of the symmetric vacuum and restores cluster decomposition [^1] Violations of cluster decomposition, aka infinite range correlations, are generally associated with spontaneous symmetry breaking..
@@
    In the unstable $\cal{T}$}-symmetric vacuum correlation functions do not match semiclassical expectations; for example two-point functions are necessarily real in the symmetric vacuum but are complex in the semiclassical theory. In the stable asymmetric vacuum correlation functions are consistent with semiclassical expectations.
@@ 2.1 The System
The subscript $s$ means system.
@@ 3.1 Breakdown of Clustering
Call the spin at $p$ the “observer” and the region near $h$ the laboratory, i.e., the place where experiments take place. The laboratory may well contain a genuine observer (no quotes) but that is incidental.
@@
Restoring cluster decomposition can be done in a number of equivalent ways, all of which pick out the “observer” as a reference point to which other spins can be compared (dressed). The first way is to introduce an addition to the Hamiltonian which only involves $p.$
@@
The symmetric ground state (3.10) is unstable to a tiny perturbation which not only polarizes the “observer” spin but the entire lattice. For $N\to\infty$ the magnitude of the effect becomes independent of the magnitude of $b.$
@@ 3.2 Gauging ${\cal{Z}}_{2}$
Adding an extra symmetry breaking as in (3.16) is artificial device which has no analog in de Sitter space. Another way to accomplish the same result—cluster decomposition—is to gauge gauge ${\cal{Z}}_{2}$ by restricting the physical space of states and the observables to be ${\cal{Z}}_{2}$ invariant. As before we single out $p$ as the “observer” and “dress” all observables to $p.$ This simply means replace all other  [^5] In this context the index $i$ runs over all spins other than the observer $Z_{p}.$ $Z_{i}$ by the dressed operators $\bar{Z}_{i},$
@@ 4.2 Clocks in the Static Patch
One feature of SSB is that the instability leading to a properly clustering state can be kicked off by a very weak perturbation. In fact almost any perturbation will work and lead to the same global behavior. This meshes well with the literature in which the “observer” can be almost anything from a galaxy to a few low energy photons.
@@ 4.3 The Puzzle
Physical clocks and spontaneous breaking of $\cal{T}$} can potentially resolve the paradox of [7]. Simply stated, the existence of a tiny time-oriented feature, arbitrarily far from the laboratory (near the horizon) will set off an instability that leads to a non-zero expectation value for $C,$ consistent with the non-zero value in the semiclassical theory.
5 Gauging $CT$ and Dressing C
@@
which rendered them gauge invariant and properly clustering. While $Z_{i}$ is odd under $K$ and is therefore not gauge invariant, $\bar{Z}_{i}$ is invariant.
@@ 6 Another Viewpoint
It is interesting to consider the flat-space limit in which $\ell\to\infty$ and the pode moves infinitely far from the horizon. One can take the flat-space limit in two ways. One way is to focus attention on the near horizon region which includes everything whose proper distance from the horizon remains finite in the limit $\ell\to\infty.$ This is the limit in which the static patch tends to Rindler space and the pode disappears to infinity.
@@
So far I have thought of the “observer” as living at the pode and the laboratory including $C$ as being near (but not too near) the horizon. This seems strange because one would expect that when the observer moves off to infinity its presence should become irrelevant to the physics in the laboratory. The resolution of this puzzle—SSB—is what this paper is all about. But it is also interesting to center the the laboratory at the pode and let the “observer” follow the horizon (or stretched horizon).
In this latter viewpoint the clock is a tiny out-of-equilibrium feature near the horizon (in the Hawking-Unruh radiation) as long as it breaks time-reversal symmetry. (Breaking $\cal{T}$} is very generic; not breaking it requires fine-tuning.)
@@ 7 Conclusion
but for the last case where the energy of the feature is $\sim T_{gh}$ the probability is order unity. One can summarize this with the phrase, “Observers” are common; Observers are rare.
In fact it is probably true that the theory without an “observer” is inconsistent; a state with a feeble observer is not orthogonal to the maximally mixed state. Projecting onto a state in which every conceivable “observer” is absent probably leaves nothing. But states without observers are consistent. If for example an observer means anything with mass greater than the Planck scale the probability to not have an observer is close to one,
@@
By contrast, the probability to not have an “observer” of any kind may be zero.
Finally we can ask about stability of the vacuum: does it require “observers” or to put it another way is the Hartle-Hawking vacuum unstable. Its holographic dual—the maximally mixed state is unstable. The answer, based on Euclidean gravitational path integration is yes, the HH state is unstable. The argument was originally due to Polchinski [8]. Recently it was revisited in [3] and [4]. The conclusion of these later papers is that the presence of an observer in the path integral is sufficient to stabilize the vacuum. It is almost certainly true that a feeble “observer” will also suffice and I will assume that this is the case.
Whether the Polchinski instability is related to the SSB instability discussed in this paper is an open question. Both instabilities are eliminated by the introduction of an “observer.”
@@
Finally a question: If $\cal{T}$} is a spontaneously broken discrete gauge symmetry (Higgs phenomenon) are there consequences of this? In particular does parallel transport around the bifurcate horizon induce a $\cal{T}$} reversal?

### v2

@@
It is widely thought that the quantum theory of de Sitter space requires the existence of a physical observer in the static patch. What exactly is meant by an observer is unclear; it could be anything from a few photons with energy just above the Gibbons-Hawking temperature to a gravitationally bound cluster of galaxies. In a recent note I explained that one way the need for observers can arise is from the spontaneous breaking of time-reversal symmetry. This longer paper expands on the subject, filling in conceptual arguments that were implicit but not explicitly stated in the earlier paper.
@@ 1 Introduction
There is a widespread belief that the quantum mechanics of de Sitter space only makes sense if there is an observer present [1]. Discrepancies between semiclassical gravity and static patch holography have been resolved by introducing a “feature” localized at the pode of the static patch. According to the terminology of [1][2][3][4] this feature is called an observer, although it may have little to do with conventional observation by almost classical apparatuses.
@@
In this paper I will be concerned with one particular discrepancy. According to static patch holography two-point correlation functions in de Sitter space are real; the imaginary parts vanish [6]. This is in contrast to bulk field theory in a de Sitter background where the two-point functions are complex. The ability of a observer, localized at the pode, to resolve this kind of discrepancy is surprising because the semiclassical calculation may involve bulk degrees of freedom localized extremely far from the pode—for example near the horizon—and yet no matter how far away the pode is from the region of the semiclassical calculation, the presence of an observer at the pode changes things enough to resolve the discrepancy. On the face of it this is a form of non-locality forbidden in bulk quantum field theory.
@@
Generally the literature on this subject does not spell out exactly what is meant by an observer. In fact it seems that it could be anything from a galaxy to a tiny out-of-equilibrium feature in the Hawking radiation—no more than a handful of photons with energy of order the Gibbons-Hawking temperature. Calling such a feeble phenomenon an observer seems misleading to me; I prefer feature to the overly suggestive observer. It is important to keep in mind that feature is just a name for an out-of-equilibrium fluctuation that could be almost anything.
Features comprise a more general class than observers. Observers, among other requirements, are heavy enough to be localized near the pode; Features can be heavy, i.e., they can be observers, or they can much lighter and can have wavelengths of cosmic scale. Localized observers will be useful for arguing that there is a breakdown of cluster-decomposition in de Sitter space, but almost any kind of feature is sufficient to resolve the mismatch.
One thing about observers or features is that they should have some form of time-reversal asymmetry to distinguish future and past. I’ll call such an asymmetric feature a “clock.”
@@
Now let’s turn to the region $h$ in figure 1. I’ll refer to it as the “laboratory.” It is the region of interest where observable phenomena take place, the phenomena being described by a set of correlation functions localized near $h.$
@@
In a recent note [5] I argued that the need for an observer (or feature) has nothing to do with the consistency of quantum mechanics; it is due to a spontaneous breaking (SSB) of time-reversal symmetry. The feature kicks off an instability of the symmetric vacuum and restores cluster decomposition [^1] Violations of cluster decomposition, aka infinite range correlations, are generally associated with spontaneous symmetry breaking..
@@
    In the unstable $\cal{T}$}-symmetric vacuum correlation functions do not match semiclassical expectations; for example two-point functions are necessarily real in the symmetric vacuum but are complex in the semiclassical theory [6]. In the stable asymmetric vacuum correlation functions are consistent with semiclassical expectations.
@@ 2.1 The System
The subscript $\sigma$ means system.
@@ 3.1 Breakdown of Clustering
Call the spin at $p$ the “observer” and the region near $h$ the “laboratory,” i.e., the place where experiments take place. The laboratory may well contain a genuine observer but that is incidental.
@@
Restoring cluster decomposition can be done in a number of equivalent ways, all of which pick out the observer as a reference point to which other spins can be compared (dressed). The first way is to introduce an addition to the Hamiltonian which only involves $p.$
@@
The symmetric ground state (3.10) is unstable to a tiny perturbation which not only polarizes the observer spin but the entire lattice. For $N\to\infty$ the magnitude of the effect becomes independent of the magnitude of $b.$
@@ 3.2 Gauging ${\cal{Z}}_{2}$
Adding an extra symmetry breaking as in (3.16) is artificial device which has no analog in de Sitter space. Another way to accomplish the same result—cluster decomposition—is to gauge gauge ${\cal{Z}}_{2}$ by restricting the physical space of states and the observables to be ${\cal{Z}}_{2}$ invariant. As before we single out $p$ as the observer and “dress” all observables to $p.$ This simply means replace all other  [^5] In this context the index $i$ runs over all spins other than the observer $Z_{p}.$ $Z_{i}$ by the dressed operators $\bar{Z}_{i},$
@@ 4.2 Clocks in the Static Patch
One feature of SSB is that the instability leading to a properly clustering state can be kicked off by a very weak perturbation. In fact almost any perturbation will work and lead to the same global behavior. This meshes well with the literature in which the observer can be almost anything from a galaxy to a few low energy photons.
@@ 4.3 The Puzzle
Physical clocks and spontaneous breaking of $\cal{T}$} can potentially resolve the paradox of [6]: the existence of a tiny time-oriented feature arbitrarily far from the laboratory (near the horizon) will set off an instability that leads to a non-zero expectation value for $C,$ consistent with the semiclassical theory.
5 Gauging $\cal{T}$} and Dressing C
@@
which rendered them gauge invariant and properly clustering. While $Z_{i}$ is odd under $K$ and is therefore not gauge invariant, $\bar{Z}_{i}$ is gauge invariant.
@@ 6 Another Viewpoint
It is interesting to consider the flat-space limit in which the de Sitter radius $\ell\to\infty,$ and the pode moves infinitely far from the horizon. One can take the flat-space limit in two ways. One way is to focus attention on the near horizon region which includes everything whose proper distance from the horizon remains finite in the limit $\ell\to\infty.$ This is the limit in which the static patch tends to Rindler space and the pode disappears to infinity.
@@
So far I have thought of the observer as living at the pode and the laboratory including $C$ as being near (but not too near) the horizon. This seems strange because one would expect that when the observer moves off to infinity its presence should become irrelevant to the physics in the laboratory. The resolution of this puzzle—SSB—is what this paper is all about. But it is also interesting to center the the laboratory at the pode and let the observer follow the horizon (or stretched horizon).
In this latter viewpoint the clock is a tiny out-of-equilibrium feature near the horizon as long as it breaks time-reversal symmetry. (Breaking $\cal{T}$} is very generic; not breaking it requires fine-tuning.)
@@ 7 Conclusion
but for the last case where the energy of the feature is $\sim T_{gh}$ the probability is order unity. One can summarize this with the phrase: Features are common; Observers are rare.
In fact it is probably true that the theory without a feature is inconsistent; a state with a feeble feature is not orthogonal to the maximally mixed state. Projecting onto a state in which every conceivable feature is absent probably leaves nothing. But states without observers are consistent. If for example an observer means anything with mass greater than the Planck scale the probability to not have an observer is close to one,
@@
By contrast, the probability to not have an feature of any kind may be zero.
Finally we can ask about stability of the vacuum: does it require observers or to put it another way is the Hartle-Hawking vacuum unstable. Its holographic dual—the maximally mixed state is unstable. The answer, based on Euclidean gravitational path integration is yes, the HH state is unstable. The argument was originally due to Polchinski [8]. Recently it was revisited in [3] and [4]. The conclusion of these later papers is that the presence of an observer in the path integral is sufficient to stabilize the vacuum. It is likely that a feeble feature will also suffice.
Whether the Polchinski instability is related in any way to the SSB instability discussed in this paper is an open question. Both instabilities are eliminated by the introduction of an observer but the Polchinski instability does not obviously have anything to do with time-reversal.